Find the for e between the two charge particles q1= +5 micro-Coulombs and q2= -1 micro-Coulombs and separated 5cm


1
Expert's answer
2021-03-21T11:30:58-0400

Given,

"q_1=5\mu C"

"q_2=-1\mu C"

"d=5cm"

"F=\frac{q_1\times q_2}{4\pi \epsilon d^2}"


"=\frac{5\times 10^{-12}\times 9\times 10^{9}}{(0.05)^2}"

"=18N"